Larry has a mass of 90 kg and runs across in front of the classroom at a speed of 5MS and jumps into a giant skateboard skateboard initially is at rest and has a mass equal to Larrys with Nore friction with the floor what is the speed of leery in the skateboard
Answer #1

using conservation of momentum

m1 v1 = (m1 + m2l v

90* 5 = ( 90 + 90) v

v = 2.5 m/s

=======

Comment before rate in case any doubt, will reply for sure.. goodluck